OpenStudy (anonymous):

Evaluate the expression below when x = 3. -2x - 8 -(-3x + 8)

OpenStudy (anonymous):

Firstly solve the brackets : Multiply - to brackets first...

OpenStudy (anonymous):

Can you do that??

OpenStudy (anonymous):

-1?

OpenStudy (anonymous):

Yes distribute -1 to the brackets... Don't plug in the value x = 3 now...

OpenStudy (anonymous):

your confusing me

OpenStudy (anonymous):

I think you have substituted the value : According to you : (-3(3) + 8) = (-9 + 8) = -1 Right ??

OpenStudy (anonymous):

So now you have got : -2x - 8 - (-1) Right ??

OpenStudy (anonymous):

okay so how do i solve that? do i plug in the 1

OpenStudy (anonymous):

No.. again substitute x = 3 in place of x.. And tell me what did you get after that..

OpenStudy (anonymous):

-10

OpenStudy (anonymous):

Can you show how you got this ??

OpenStudy (anonymous):

i plugged in the 3 for x.

OpenStudy (anonymous):

Can you show me that??

OpenStudy (anonymous):

..............................

OpenStudy (anonymous):

Ha ha ha ha... Try man, you can do that.. See : you have -2x - 8 here put x = 3 and tell me what did you get ??